You are planning to make monthly deposits of $170 into a retirement account that pays 13 percent interest compounded monthly. If your first deposit will be made one month from now, how large will your retirement account be in 29 years?
$683,947.36

$618,809.51

$7,816,541.20

$527,510.09

$651,378.43
Answer #1

170 170 170 70 - 348 348 Using financial calculators N = 348 PMT= 170 1/4 = 134 = 1.084/ CPT fv $651378-43 $651378.43 .: Valu
Answer #2

SOLUTION :


Monthly compounding.


Monthly deposit, A = 170 ($)


Monthly interest rate, r = 13/12 % = 13/1200 in fractions.

=> 1 + r = 1213/1200 


Periods, n = 29 * 12 = 348 months


So, 


FV after 29 years (348 months)


= A((1 + r)^n - 1) / r

= 170 * ((1213/1200)^348 - 1) / (13/1200)

= 651378.43 ($)  : 4th OPTION (ANSWER).

answered by: Tulsiram Garg